OpenStudy (anonymous):

determine whether the system has one solution, no solution, or indefinitely many solutions

OpenStudy (anonymous):

y=2/3x 2x-y=-4

OpenStudy (lgbasallote):

\(\large y = \frac{2}{3}x?\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Solve one equation for one of the two variables. In this case, the first equation is already solved for y. Substitute in for that variable in the second equation, simplify, and see what you get.

OpenStudy (anonymous):

Is it an equation that is always true, like 1=1? Is it an equation that is never true, like 1=2? Or is it an equation that is true for one value of x, like x=3?

OpenStudy (anonymous):

Try it out and ask me any questions that come up.

OpenStudy (precal):

 if your two lines cross, then you will have one solution